OpenStudy (anonymous):

Find the vertex of the parabola and give its direction of opening. x=1/2y^2+3y-6?

OpenStudy (anonymous):

Can someone just tell me if my answer is right? The vertex is (-21/2, -3). <---- *opens to the right*

OpenStudy (anonymous):

@timo86m is that correct

OpenStudy (anonymous):

you can graph it

OpenStudy (anonymous):

So it opens up?!!

OpenStudy (anonymous):

it is odd could be a trick questions since it is f(y) not f(x) lol

OpenStudy (anonymous):

how'd you do it did you solve for y?

OpenStudy (anonymous):

x= 1/2y^2 + 3y - 6 x = (1/2)(y^2 + 6y) - 6 x = (1/2)(y^2 + 6y + 9) - 6 - 9/2 x = (1/2)(y + 3)^2 - 21/2. The vertex is (-21/2, -3). <---- *opens up*

OpenStudy (anonymous):

Thats correct

OpenStudy (anonymous):

It opens to the right, never mind

OpenStudy (anonymous):

i just graphed it

OpenStudy (anonymous):

that is what i said solve for y it is in f(y) form not f(x) form

OpenStudy (anonymous):